### v4

@@ Appendix A Auxiliary Results
$\lim_{\ell\uparrow\infty}{h_{\ell}^{-\frac{1}{2}}\mathbb{E}[(e^{h_{\ell}}_{1}-e^{h_{\ell-1}}_{1})^{2}]}=\frac{1}{(1-\alpha)^{2}}\lim_{\ell\uparrow\infty}{h_{\ell}^{-\frac{1}{2}}\mathbb{E}\big[\big|\mathds{1}_{X_{h_{\ell}}>\xi^{0}_{\star}}-\mathds{1}_{X_{h_{\ell-1}}>\xi^{0}_{\star}}\big|\big]}=\frac{\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)^{2}}.$ (86)
On the other hand, reusing the notation in Step 7.1 of Theorem 1.2’s proof, Appendix B, for $\varepsilon>0$, there exists $n_{0}\geq 0$ such that, for $n\geq n_{0}$, $1-(\mu+\varepsilon)\gamma_{n}>0$ and $|\Delta\Sigma_{n}|\leq\varepsilon$. There also exists $L_{0}\geq 1$ such that, for $L\geq L_{0}$, one has $N_{1}\geq\dots\geq N_{L}\geq n_{0}$, so that $|\Delta\Sigma_{N_{L}}|\leq\varepsilon$. Let $L\geq L_{0}$. By (56) and Lemma A.1(i), for $n\geq N_{L}$,
@@
$\limsup_{L\uparrow\infty}\sup_{1\leq\ell\leq L}|\Delta\Sigma_{N_{\ell}}|=0.$ (87)
To sum up, using (86) and (87), Cesàro’s lemma ($\infty$’ version) yields
@@
Thus, reusing (86), by Cesàro’s lemma ($\infty$’ version),
@@
$=\frac{h_{L}^{-\frac{2}{\beta}-\frac{2\beta-1}{2\beta(1+\beta)}}}{(1-\alpha)^{2}}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big).$
By the $1$-Lipschitz property of $\mapstochar\rightarrow x^{+}, xx\in\mathbb{R}$, and Lemma A.3(i),
@@
$h_{\ell}^{-\frac{1}{2}}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits_{\ell\uparrow\infty}^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G.$ (88)
Besides, the $1$-Lipschitz property of $\mapstochar\rightarrow x^{+}, xx\in\mathbb{R}$, and Lemmas A.2(i) and A.3(i) guarantee that, for $2<p\leq 2+\delta$,
$\sup_{\ell\geq 1}h_{\ell}^{-\frac{p}{2}}\,\mathbb{E}\big[\big|(X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big|^{p}\big]<\infty.$ (89)
Eventually, the weak convergence (88) and the uniform integrability (89) yield
$\lim_{\ell\uparrow\infty}{h_{\ell}^{-1}\,\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)}=\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G).$ (90)
Ultimately, via (21) and Cesàro’s lemma (‘$0/0$’ version),
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{2\beta-1}{2(1+\beta)}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{2\beta-1}{2(1+\beta)}}h_{\ell}^{-1}\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
$=\frac{h_{0}^{\frac{2\beta-1}{2(1+\beta)}}\big(M^{\frac{2\beta-1}{2(1+\beta)}}-1\big)^{\frac{1}{\beta}-1}}{(1-\alpha)^{2}}\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G).$
@@
We prove hereby that $\big\{S^{1,2}_{L},L\geq 1\big\}$ converges to 0 in $L^{1}(\mathbb{P})$. It follows from (86) and (90) that
$\mathbb{E}[(e^{h_{\ell}}_{1}-e^{h_{\ell-1}}_{1})^{2}]$ $=\frac{1}{(1-\alpha)^{2}}\,\mathbb{E}\big[\big|\mathds{1}_{X_{h_{\ell}}>\xi^{0}_{\star}}-\mathds{1}_{X_{h_{\ell-1}}>\xi^{0}_{\star}}\big|\big]\leq Ch_{\ell}^{\frac{1}{2}},$ (91)
$\mathbb{E}[(\eta^{h_{\ell}}_{1}-\eta^{h_{\ell-1}}_{1})^{2}]$ $=\frac{1}{(1-\alpha)^{2}}\mathbb{V}\mathrm{ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\leq Ch_{\ell},$
@@
$\mathbb{E}\big[\big|(e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k})(\eta^{h_{\ell}}_{k}-\eta^{h_{\ell-1}}_{k})\big|\big]\leq Ch_{\ell}^{\frac{3}{4}}.$ (92)
Using (92), (45), the fact $\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>0$ by Remark 1.1(ii), Lemma A.1(i) and (21), we obtain
@@
$S^{1,2}_{L}=S^{2,1}_{L}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}\text{\rm-as}}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\widetilde{\Sigma}_{\beta}\coloneqq\begin{pmatrix}\frac{\gamma_{1}\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)(2f_{X_{0}}(\xi^{0}_{\star})-(1-\alpha)\gamma_{1}^{-1}\mathds{1}_{\beta=1})}&0\\ 0&\frac{h_{0}^{\frac{2\beta-1}{2(1+\beta)}}\big(M^{\frac{2\beta-1}{2(1+\beta)}}-1\big)^{\frac{1}{\beta}}}{(1-\alpha)^{2}}\frac{\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi\star}\,G)}{M^{\frac{2\beta-1}{2(1+\beta)}}-1}\end{pmatrix}$.
@@
Recalling the definitions (63) and (36)–(39), the decomposition (35) yields
$\xi^{h}_{n}-\xi^{h}_{\star}=\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}a^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}g^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}r^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}\rho^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}e^{h}_{n}.$ (93)
@@
$\quad+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k}.$ (94)
@@
Owing to (13), Lemma A.3(ii), a comparison between series and integrals and the fact that $\beta<1$,
@@
$h_{L}^{-1}\big(\overline{\xi}^{h_{0}}_{N_{0}}-\xi^{h_{0}}_{\star}\big)\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}a^{h_{\ell}}_{k}-a^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Using (79) and a comparison between series and integrals, we obtain
@@
If $\delta_{0}\geq\frac{3}{4}$, by (27) and the fact that $\beta<1$,
@@
Otherwise, if $\frac{1}{8}\leq\delta_{0}<\frac{3}{4}$, reusing (27),
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}g^{h_{\ell}}_{k}-g^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Using (48), a comparison between series and integrals and (27), we obtain
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}r^{h_{\ell}}_{k}-r^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Reusing similar arguments to Step 5 of the proof of Theorem 3.5, Appendix E, we obtain via (49), (27) and a comparison between series and integrals,
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k},\ell\geq 1\big\}$ are independent and centered. Hence, using (50), a comparison between series and integrals and (27),
@@
$h_{L}^{-\frac{9}{8}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$h_{L}^{-\frac{9}{8}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\zeta^{h_{\ell}}_{k}-\zeta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
It follows from (27) that
@@
From (50) and (51), we get
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\theta^{h_{0}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad N_{0}\uparrow\infty,$
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\zeta^{h_{0}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad N_{0}\uparrow\infty.$
Since $\{\eta^{h_{0}}_{k},k\geq 1\}$ is i.i.d. such that $\mathbb{E}[|\eta^{h_{0}}_{1}|^{2}]<\infty$ by (44) and the fact that, by assumption, $\mathbb{E}[|\varphi(Y,Z)|^{2+\delta}]<\infty$, by the classical CLT,
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\eta^{h_{0}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\Big(0,\frac{\mathbb{V}\mathrm{ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\Big)\quad\mbox{ as }\quad N_{0}\uparrow\infty.$
@@
$h_{L}^{-\frac{9}{8}}\big(\chi^{h_{0}}_{N_{0}}-\chi^{h_{0}}_{\star}\big)\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\Big(0,h_{0}^{-\frac{3}{4}}\big(1-M^{-\frac{1}{4}}\big)\frac{\mathbb{V}\mathrm{ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\Big)\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Since, for all $k\geq 1$, $\big\{\frac{1}{N_{\ell}}(e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k})\mathds{1}_{1\leq k\leq N_{\ell}},1\leq\ell\leq L\big\}$ are independent and centered, using the Marcinkiewicz-Zygmund inequality, Jensen’s inequality and (83),
@@
Proceeding as in Step 9.1.2. in Theorem 3.5’s proof, Appendix E, using in particular (85), that $\sup_{h\in\mathcal{H}}\mathbb{E}[|X_{h}|^{2+\delta}]<\infty$ by assumption and (27),
@@
By (27), (86) and Cesàro’s lemma ($\infty$’ version),

### v5

@@ Appendix A Auxiliary Results
$\lim_{\ell\uparrow\infty}{h_{\ell}^{-\frac{1}{2}}\mathbb{E}[(e^{h_{\ell}}_{1}-e^{h_{\ell-1}}_{1})^{2}]}=\frac{1}{(1-\alpha)^{2}}\lim_{\ell\uparrow\infty}{h_{\ell}^{-\frac{1}{2}}\mathbb{E}\big[\big|\mathds{1}_{X_{h_{\ell}}>\xi^{0}_{\star}}-\mathds{1}_{X_{h_{\ell-1}}>\xi^{0}_{\star}}\big|\big]}=\frac{\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)^{2}}.$ (E.10)
On the other hand, reusing the notation in Step 7.1 of Theorem 1.4’s proof, Appendix B, for $\varepsilon>0$, there exists $n_{0}\geq 0$ such that, for $n\geq n_{0}$, $1-(\mu+\varepsilon)\gamma_{n}>0$ and $|\Delta\Sigma_{n}|\leq\varepsilon$. There also exists $L_{0}\geq 1$ such that, for $L\geq L_{0}$, one has $N_{1}\geq\dots\geq N_{L}\geq n_{0}$, so that $|\Delta\Sigma_{N_{L}}|\leq\varepsilon$. Let $L\geq L_{0}$. By (B.22) and Lemma A.1(i), for $n\geq N_{L}$,
@@
$\limsup_{L\uparrow\infty}\sup_{1\leq\ell\leq L}|\Delta\Sigma_{N_{\ell}}|=0.$ (E.11)
To sum up, using (E.10) and (E.11), Cesàro’s lemma ($\infty$’ version) yields
@@
Thus, reusing (E.10), by Cesàro’s lemma ($\infty$’ version),
@@
$=\frac{h_{L}^{-\frac{2}{\beta}-\frac{2\beta-1}{2\beta(1+\beta)}}}{(1-\alpha)^{2}}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big).$
By the $1$-Lipschitz property of $\mapsto x^{+}, xx\in\mathbb{R}$, and Lemma A.3(i),
@@
$h_{\ell}^{-\frac{1}{2}}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\mathrel{{\mathop{\longrightarrow}\limits_{\ell\uparrow\infty}^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G.$ (E.12)
Besides, the $1$-Lipschitz property of $\mapsto x^{+}, xx\in\mathbb{R}$, and Lemmas A.2(i) and A.3(i) guarantee that, for $2<p\leq 2+\delta$,
$\sup_{\ell\geq 1}h_{\ell}^{-\frac{p}{2}}\,\mathbb{E}\big[\big|(X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big|^{p}\big]<\infty.$ (E.13)
Eventually, the weak convergence (E.12) and the uniform integrability (E.13) yield
$\lim_{\ell\uparrow\infty}{h_{\ell}^{-1}\,\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)}=\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G).$ (E.14)
Ultimately, via (3.7) and Cesàro’s lemma (‘$0/0$’ version),
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}\frac{h_{\ell}}{N_{\ell}}h_{\ell}^{-1}\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
@@
$\quad\times\lim_{L\uparrow\infty}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{2\beta-1}{2(1+\beta)}}\bigg)^{-1}\bigg(\sum_{\ell=1}^{L}h_{\ell}^{\frac{2\beta-1}{2(1+\beta)}}h_{\ell}^{-1}\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\bigg)$
$=\frac{h_{0}^{\frac{2\beta-1}{2(1+\beta)}}\big(M^{\frac{2\beta-1}{2(1+\beta)}}-1\big)^{\frac{1}{\beta}-1}}{(1-\alpha)^{2}}\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G).$
@@
We prove hereby that $\big\{S^{1,2}_{L},L\geq 1\big\}$ converges to 0 in $L^{1}(\mathbb{P})$. It follows from (E.10) and (E.14) that
$\mathbb{E}[(e^{h_{\ell}}_{1}-e^{h_{\ell-1}}_{1})^{2}]$ $=\frac{1}{(1-\alpha)^{2}}\,\mathbb{E}\big[\big|\mathds{1}_{X_{h_{\ell}}>\xi^{0}_{\star}}-\mathds{1}_{X_{h_{\ell-1}}>\xi^{0}_{\star}}\big|\big]\leq Ch_{\ell}^{\frac{1}{2}},$ (E.15)
$\mathbb{E}[(\eta^{h_{\ell}}_{1}-\eta^{h_{\ell-1}}_{1})^{2}]$ $=\frac{1}{(1-\alpha)^{2}}\operatorname{\mathbb{V}ar}\big((X_{h_{\ell}}-\xi^{h_{\ell}}_{\star})^{+}-(X_{h_{\ell-1}}-\xi^{h_{\ell-1}}_{\star})^{+}\big)\leq Ch_{\ell},$
@@
$\mathbb{E}\big[\big|(e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k})(\eta^{h_{\ell}}_{k}-\eta^{h_{\ell-1}}_{k})\big|\big]\leq Ch_{\ell}^{\frac{3}{4}}.$ (E.16)
Using (E.16), (B.11), the fact $\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>0$ by Remark 1.2(ii), Lemma A.1(i) and (3.7), we obtain
@@
$S^{1,2}_{L}=S^{2,1}_{L}\mathrel{{\mathop{\longrightarrow}\limits^{\operatorname{\mathds{P}\text{-as}}}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\widetilde{\Sigma}_{\beta}\coloneqq\begin{pmatrix}\frac{\gamma_{1}\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)(2f_{X_{0}}(\xi^{0}_{\star})-(1-\alpha)\gamma_{1}^{-1}\mathds{1}_{\beta=1})}&0\\ 0&\frac{h_{0}^{\frac{2\beta-1}{2(1+\beta)}}\big(M^{\frac{2\beta-1}{2(1+\beta)}}-1\big)^{\frac{1}{\beta}}}{(1-\alpha)^{2}}\frac{\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi\star}\,G)}{M^{\frac{2\beta-1}{2(1+\beta)}}-1}\end{pmatrix}$.
@@
Recalling the definitions (C.2) and (B.2)–(B.5), the decomposition (B.1) yields
$\xi^{h}_{n}-\xi^{h}_{\star}=\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}a^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}g^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}r^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}\rho^{h}_{n}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}e^{h}_{n}.$ (F.1)
@@
$\quad+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}+\frac{1}{V_{0}^{\prime\prime}(\xi^{0}_{\star})}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k}.$ (F.2)
@@
Owing to (2.1), Lemma A.3(ii), a comparison between series and integrals and the fact that $\beta<1$,
@@
$h_{L}^{-1}\big(\overline{\xi}^{h_{0}}_{N_{0}}-\xi^{h_{0}}_{\star}\big)\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}a^{h_{\ell}}_{k}-a^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Using (E.3) and a comparison between series and integrals, we obtain
@@
If $\delta_{0}\geq\frac{3}{4}$, by (4.2) and the fact that $\beta<1$,
@@
Otherwise, if $\frac{1}{8}\leq\delta_{0}<\frac{3}{4}$, reusing (4.2),
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}g^{h_{\ell}}_{k}-g^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Using (B.14), a comparison between series and integrals and (4.2), we obtain
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}r^{h_{\ell}}_{k}-r^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Reusing similar arguments to Step 5 of the proof of Theorem 3.5, Appendix E, we obtain via (B.15), (4.2) and a comparison between series and integrals,
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k},\ell\geq 1\big\}$ are independent and centered. Hence, using (B.16), a comparison between series and integrals and (4.2),
@@
$h_{L}^{-\frac{9}{8}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$h_{L}^{-\frac{9}{8}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\zeta^{h_{\ell}}_{k}-\zeta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
It follows from (4.2) that
@@
From (B.16) and (B.17), we get
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\theta^{h_{0}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad N_{0}\uparrow\infty,$
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\zeta^{h_{0}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad N_{0}\uparrow\infty.$
Since $\{\eta^{h_{0}}_{k},k\geq 1\}$ is i.i.d. such that $\mathbb{E}[|\eta^{h_{0}}_{1}|^{2}]<\infty$ by (B.10) and the fact that, by assumption, $\mathbb{E}[|\varphi(Y,Z)|^{2+\delta}]<\infty$, by the classical CLT,
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\eta^{h_{0}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\Big(0,\frac{\operatorname{\mathbb{V}ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\Big)\quad\mbox{ as }\quad N_{0}\uparrow\infty.$
@@
$h_{L}^{-\frac{9}{8}}\big(\chi^{h_{0}}_{N_{0}}-\chi^{h_{0}}_{\star}\big)\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\Big(0,h_{0}^{-\frac{3}{4}}\big(1-M^{-\frac{1}{4}}\big)\frac{\operatorname{\mathbb{V}ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\Big)\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Since, for all $k\geq 1$, $\big\{\frac{1}{N_{\ell}}(e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k})\mathds{1}_{1\leq k\leq N_{\ell}},1\leq\ell\leq L\big\}$ are independent and centered, using the Marcinkiewicz-Zygmund inequality, Jensen’s inequality and (E.7),
@@
Proceeding as in Step 9.1.2. in Theorem 3.5’s proof, Appendix E, using in particular (E.9), that $\sup_{h\in\mathcal{H}}\mathbb{E}[|X_{h}|^{2+\delta}]<\infty$ by assumption and (4.2),
@@
By (4.2), (E.10) and Cesàro’s lemma ($\infty$’ version),